Gwen mixes 5 pounds of dried cranberries that sell for $4 per pound with some dried apples that sell for $3 per pound. If she wants the mix to be worth $3.25 per pound, then how many pounds of apples should she use?
Answer by mananth(16946)

Gwen mixes 5 pounds of dried cranberries that sell for $4 per pound with some dried apples that sell for $3 per pound. If she wants the mix to be worth $3.25 per pound, then how many pounds of apples should she use?
..
apples be x lbs.
cost = $3
..
cranberries 5 lbs
cost $4
..
Mix x+5 lbs
cost $ 3.25
..
3x+4*5=3.25(x+5)
3x+20=3.25x+16.25
0.25x=20-16.25
0.25x=3.75
x= 3.75/0.25
x=15 lbs of apples